OpenStudy (anonymous):

find the standard equation of a line with slope 3/4 and y-intercept of -1

OpenStudy (anonymous):

\[y=mx+b\] \[m=\frac{3}{4},b--1\]

OpenStudy (anonymous):

A. 4X-3Y=3 B. 4Y-3X=4 C. 3X+4Y=-4 D. 3X-4Y=4

OpenStudy (anonymous):

D.